Let h and j be function whose first and second derivatives exist on interval I. Which of the following formula(s) is true. Explain.
a. f g" - f" g = (f g' - f' g)'
b. f g" + f" g= (fg)"

Answer by robertb(5830)

a.is true, because (fg'-f'g)' = f'g'+fg" - (f"g+f'g') = fg"-f"g.
b. is false, because (fg)" = (f'g +fg')' = f"g + f'g' + f'g' + fg" = f"g +
2f'g' + fg"